OpenStudy (anonymous):

Show that f'(x)=0 at some point between the two x-intercepts of the function f(x) = -6x√x+1 x intercepts (-1,0) and (0,0) Find a value of x such that f'(x) = 0 x=

OpenStudy (anonymous):

This is calc 1 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Not really sure about the first part but the way you find value for x such that f'(x)=0 by taking the derivative of the given function and then setting it to zero and solving for x

OpenStudy (anonymous):

I'm thinking for the first part you can plug in the intercepts and take the derivative of them to show that f;(x) is in between

OpenStudy (anonymous):

\[-6x \sqrt{x+1} = -6x[\frac{ 1 }{ 2 }(x+1)]^{-1/2}\]

OpenStudy (anonymous):

Actually there is only one part, finding the value of x

OpenStudy (anonymous):

Well it is asking you to show something, and then it asks you to find something.

OpenStudy (anonymous):

I think they mean the same thing

OpenStudy (anonymous):

Is this a problem in a book?

OpenStudy (anonymous):

From an online course. The original question was: Find the two x-intercepts of the function f and show that f '(x) = 0 at some point between the two x-intercepts. The second part of the question was: Find a value of x such that f '(x) = 0.

OpenStudy (anonymous):

Okay, for the second part I believe you take the derivative of f(x) and then set f'(x)=0. Then solve for x.

OpenStudy (gorv):

\[x*\sqrt{x}=x^{3/2}\]

OpenStudy (gorv):

\[f'(x)=-6*\frac{ 3 }{ 2 }*x^{1/2}+0\]

OpenStudy (gorv):

f'(x)=0=-9sqrtx=0

OpenStudy (gorv):

x=0

OpenStudy (anonymous):

I got x=-2/3

OpenStudy (anonymous):

\[f'(x)=-6\sqrt{x+1}-\frac{ 3x }{ \sqrt{x+1} }\]

OpenStudy (anonymous):

by the product rule

OpenStudy (anonymous):

I'm confused at where that 3 comes from

OpenStudy (anonymous):

Has your teacher covered the product rule yet?

OpenStudy (anonymous):

yes. we've been through that

OpenStudy (anonymous):

Derivative of the first times the second plus the first times the derivative of the second

OpenStudy (anonymous):

actually it's an all online course so I have to learn it myself

OpenStudy (anonymous):

Essentially this: \[-6(x+1)^\frac{ 1 }{ 2 } - (6x)(\frac{ 1 }{ 2 })(x+1)^\frac{- 1 }{ 2 }\]

OpenStudy (anonymous):

I took the derivative of the first term multiplied it by the second term

OpenStudy (anonymous):

vice versa for the other half

OpenStudy (anonymous):

\[-6x(x+1)^{-1/2}(\frac{ 3 }{ 2 }x+1)=0\]

OpenStudy (anonymous):

WolframAlpha is also good

OpenStudy (anonymous):

Yeah it is.

OpenStudy (anonymous):

\[-\frac{ 3 (3x+2) }{ \sqrt{x+1} }\]

OpenStudy (anonymous):

Is that the simplified form?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

it comes from \[-6(\sqrt{1+x}+1\frac{ x }{ 2\sqrt{1+x}})\]

OpenStudy (anonymous):

I'm not sure how -2/3 comes out of that

OpenStudy (anonymous):

Let me solve real quick

OpenStudy (anonymous):

\[f'(x)=-6\sqrt{x+1}-\frac{ 3x }{ \sqrt{x+1} }\] \[0=-6\sqrt{x+1}-\frac{ 3x }{ \sqrt{x+1} }\] \[0=-6(x+1)-3x\] \[0=-6x-6-3x\] \[9x=-6\]

OpenStudy (anonymous):

\[x=-\frac{ 6 }{ 9 }=-\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

ok. think I've got it now

OpenStudy (anonymous):

Okay great. Good luck!